OpenStudy (unklerhaukus):

Solve \[y''+y'-2y=(3+4x)e^x\] for \(y=y_c+y_p\) by first finding \(y_c\) the complementry solution, and then \(y_p\) the particular integral by; i) using operator D ii) using method of undertermined coeffecients iii) using variation of parameters iv) using the fact that \(y_1(x)=e^x\) satisfies the homgenouse equation

OpenStudy (unklerhaukus):

\[\begin{align*} y''+y'-2y &=(3+4x)e^x \\ \\ y_c''+y_c'-2y_c &=0 \\ m^2+m-2 &=0 \\ (m-1)(m+2) &=0 \\ \\ y_c=Ae^x+Be^{-2x} \end{align*}\]

OpenStudy (unklerhaukus):

i) \[\newcommand \D {\text D} \begin{align*} y_p''+y_p'-2y_p &=(3+4x)e^x \\ \\ \text{Operator}\D \\ \\ (\D^2+\D-2)y_p &=(3+4x)e^x \\ y_p &=\frac1{\D^2+\D-2}(3+4x)e^x \\ &=e^x\frac1{(\D+1)^2+(\D+1)-2}(3+4x) \\ &=e^x\frac1{\D^2+2\D+1+\D+1-2}(3+4x) \\ &=e^x\frac1{\D^2+3\D}(3+4x) \\ &=e^x\frac1\D\frac1{\D+3}(3+4x) \\ &=\frac{e^x}3\frac1\D\frac1{1+\D/3}(3+4x) \\ &=\frac{e^x}3\frac1\D\Big(1-\D/3+\D^2/9-\dots\Big)(3+4x) \\%\D^3/27+ &=\frac{e^x}3\frac1\D\Big((3+4x)-4/3+0-\dots\Big) \\%0+ &=\frac{e^x}3\frac1\D\Big(\tfrac53+4x\Big) \\ &=\frac{e^x}3\int\tfrac53+4x\dd x \\ &=\frac{e^x}3\big(\tfrac53x+2x^2\big) \\%+c y_p &=\big(\tfrac59x+\tfrac23x^2\big)e^x \\ \\ y(x) &=Ae^x+Be^{-2x}+\tfrac59xe^x+\tfrac23x^2e^x \end{align*}\]

OpenStudy (unklerhaukus):

ii) \[\begin{align*} y_p''+y_p'-2y_p &=(3+4x)e^x \\ \\ \text{Method of undetermined coefficients} \\ \\ y_p =&(Dx+Ex^2)e^x \\ y'_p =&(Dx+Ex^2)e^x+(D+2Ex)e^x \\ y''_p =&(Dx+Ex^2)e^x+2(D+2Ex)e^x+(2E)e^x \\ \\ y_p''+y_p'-2y_p =&3(D+2Ex)e^x+(2E)e^x =(3+4x)e^x \\ &\qquad(3D+2E)+(6E)x=3+4x \\ \\ &\qquad\qquad\quad3D+2E =3 \\ &\qquad\qquad\qquad\qquad\quad E =\tfrac46 =\tfrac23 \\ &\qquad\qquad\qquad D =\tfrac{3-4/3}{3} =\tfrac59 \\ \\ y_p = &\left(\tfrac59x+\tfrac23x^2\right)e^x \\ \\ y = &Ae^x+Be^{-2x}+\tfrac59xe^x+\tfrac23x^2e^x \end{align*}\]

OpenStudy (unklerhaukus):

iii) \[\begin{align} y''+y'-2y &=(3+4x)e^x \\ \\ y_c''+y_c'-2y_c &=0 \\ m^2+m-2 &=0 \\ (m-1)(m+2) &=0 \\ \\ y_c=Ae^x+Be^{-2x} \end{align}\] \[\newcommand \dd [1] { \,\mathrm d#1 } \begin{align} &\text{Variation of Parameters} \\ \\ W &= \begin{vmatrix} e^x & e^{-2x} \\ e^x & -2e^{-2x} \\ \end{vmatrix} =-2e^{-x}-e^{-x} =-3e^{-x} \\ \\ y &=Ae^x+Be^{-2x} -e^x\int\frac{e^{-2x}(3+4x)e^x}{-3e^{-x}}\dd x +e^{-2x}\int\frac{e^x(3+4x)e^x}{-3e^{-x}}\dd x \\ &=Ae^x+Be^{-2x} +\frac{e^x}3\int3+4x\dd x -\frac{e^{-2x}}3\int(3+4x)e^{3x}\dd x \\ &=Ae^x+Be^{-2x}+\frac{e^x}3(3x+2x^2) -\frac{e^{-2x}}3\Big(\frac{(3+4x)e^{3x}}{3} -\int4\frac{e^{3x}}3\dd x\Big) \\ &=Ae^x+Be^{-2x}+(x+\tfrac23x^2)e^x -\frac{e^{-2x}}3\Big((1+\tfrac43x)e^{3x} -\tfrac49e^{3x}\Big) \\ &=Ae^x+Be^{-2x}+(x+\tfrac23x^2)e^x -\tfrac13\Big(\tfrac59+\tfrac43x\Big)e^{x} \\ &=Ae^x+Be^{-2x}+(x+\tfrac23x^2)e^x -\Big(\tfrac5{27}+\tfrac49x\Big)e^{x} \\ &=(A-\tfrac5{27})e^x+Be^{-2x}+(\tfrac59x+\tfrac23x^2)e^x \\ &=Ce^x+Be^{-2x}+\tfrac59xe^x+\tfrac23x^2e^x \end{align}\]

OpenStudy (unklerhaukus):

iv) \[\begin{align*} y''+y'-2y &=(3+4x)e^x \\ \\ \text{using}\qquad y_1 &=e^x \\ \\ y &=ue^x \\ y' &=\big(u'+u\big)e^x \\ y'' &=\big(u''+2u'+u\big)e^x \\ \\ y''+y'-2y &=\big(u''+3u'\big)e^x \\ &&\text{Let }u' &=p \\ && u'' &=p' \\ y''+y'-2y &=\big(3p+p'\big)e^x =(3+4x)e^x \\ p'+3p &=3+4x \\ && R=e^{\int 3\dd x}=e^{3x} \\ (pe^{3x})' &=3e^{3x}+4xe^{3x} \\ pe^{3x} &=\int3e^{3x}+4xe^{3x}\dd x \\ pe^{3x} &=e^{3x}+\tfrac43xe^{3x}-\frac43\int e^{3x}\dd x \\ pe^{3x} &=e^{3x}+\tfrac43xe^{3x}-\tfrac49e^{3x}\\ pe^{3x} &=\tfrac59e^{3x}+\tfrac43xe^{3x}\\ p &=\tfrac59+\tfrac43x \\ u' &=\tfrac59+\tfrac43x \\ u &=\int\tfrac59+\tfrac43x\dd x \\ ye^{-x} &=\tfrac59x+\tfrac23x^2 \\ y &=\tfrac59xe^x+\tfrac23x^2e^x \end{align*}\]

OpenStudy (unklerhaukus):

is there a mistake in method iii) Variation of Parameters with that constant term

OpenStudy (unklerhaukus):

should i have used y_p in iv)?

OpenStudy (abb0t):

Doesn't variantion of paremeters involve useing the wronskian?

OpenStudy (unklerhaukus):

W=|...|

OpenStudy (abb0t):

\(W_1 =\begin{bmatrix} 0 & y_1(x) & \\ f(x) & y'_1(x) & \end{bmatrix}\) and \(W_2 =\begin{bmatrix} y_2(x) & 0 & \\ y'_2(x) & f(x) & \end{bmatrix}\) to get \(u'_1 = \frac{W_1}{W}\) and \(u'_2 = \frac{W_2}{W}\) THEN INTEGRATE to get \(u_1\) and \(u_2\) or something ike that?

OpenStudy (abb0t):

I think. Im a bit rusty. so, correct me if im wrong.

OpenStudy (unklerhaukus):

Im kinda new to variation of parameters but i was using \[W(x)=\begin{vmatrix}y_1(x)&y_2(x)\\y'(x)&y_2'(x)\end{vmatrix}\] where W is the wronskina of y_1 and y_2

OpenStudy (unklerhaukus):

and \[y_p=-y_1(x)\int\frac{y_2(x)f(x)}{W(x)}\mathrm dx+y_2(x)\int\frac{y_1(x)f(x)}{W(x)}\mathrm dx\]

OpenStudy (abb0t):

Ohhhh, essentually, you are doing the same thing, but I just dont like directly going into that formula, b/c my prof's didn't like using the formula, instead they wanted to see us use matrices to get to that.

OpenStudy (unklerhaukus):

so my wronskian is your two wronskians tangled together

OpenStudy (abb0t):

Pretty much. \(W_1 =\begin{bmatrix} 0 & y_1(x) & \\ f(x) & y'_1(x) & \end{bmatrix}\) = \(0(y'_(x))-y_1(f(x)) = u_1'\) but since you want \(u_1\) then, \(\int u_1'\) and then add \(u_2'\)

OpenStudy (abb0t):

And I see you used the annihilator method for undetermined coefficients? Correct.